# A Bayesian approach to functional regression: theory and computation

### v2

@@ Appendix B More details on posterior consistency
We start by checking that $\theta\mapsto P_{\theta}(Y|X)(A_{1})$ is measurable for every measurable set $A_{1}\subseteq\mathcal{Y}$. Indeed, consider the function $F(y,\theta)=f(y|X,\theta)\bm{1}_{A_{1}}(y)$, where $f(\cdot|X,\theta)$ is the density of the normal distribution $\mathcal{N}(\alpha_{0}+\sum_{j}\beta_{j}X(t_{j}),\sigma^{2})$ and $\bm{1}_{A_{1}}$ is the indicator function of the set $A_{1}$. It is easy to see that $F(y,\theta)$ is jointly measurable (it is in fact continuous) if $X$ has continuous sample paths. Then, by Tonelli’s theorem (e.g. Folland, 1999, Theorem 2.37), the function
[…]
is measurable. Now, if $A$ is a measurable subset of $\mathcal{X}\times\mathcal{Y}$, we have the decomposition
[…]
where $A_{x}=\{y\in\mathcal{Y}:(x,y)\in A\}$ is the $x$-section of $A$. We just saw that $\theta\mapsto P_{\theta}(Y|X=x)(A_{x})$ is measurable, and thus $\theta\mapsto P_{\theta}(X,Y)(A)$ is measurable, since integration respects measurability. ∎
We also examine more closely the claim that our posterior consistency result can be refined so that the set where consistency may fail is “small” in some sense. Using the notation introduced in Section 3, and considering the Lebesgue measure on $\Theta_{p}$, denoted by $\lambda_{p}$, we need to impose an additional condition:
Condition 9 (Absolute continuity of $\lambda_{p}$).
Under the model implied by (3.3), for all $p\in\mathbb{N}$:
  1. (i)
    $\Pi(\mathcal{P}=p)>0$.
  2. (ii)
    $\sum_{\nu\in S_{p}}\Pi(\bm{\theta}[\nu]\in B|\mathcal{P}=p)=0$ implies $\lambda_{p}(B)=0$, for all $B\subseteq\Theta_{p}$ measurable.
The first condition is a somewhat technical requirement. The second condition is met, for example, if $\bm{\theta}|p$ has a density with respect to the Lebesgue measure that is invariant to permutations of the component labels and positive on all of $\Theta_{p}$. With these prerequisites in mind, denoting by $\lambda$ the extension of the Lebesgue measure to $\Theta$, we get the announced result:
Proposition 10.
Assume Condition 3 and Condition 9. Then the conclusion of Theorem 4 remains valid with $\lambda(\Theta\setminus\Theta_{*})=0$.
@@
Define $\Theta_{*}$ as in the proof of Theorem 4. Recall that $\Pi(\Theta_{*})=1$, and observe that
Since $\Pi(\mathcal{P}=p)>0$ for all $p\in\mathbb{N}$ by Condition 9-(i), we have $\Pi(\Theta_{p}\setminus\Theta_{*}|\mathcal{P}=p)=0$ for all $p$. Now, for $\nu\in S_{p}$, let $\mu^{\nu}_{p}$ be the distribution of $\bm{\theta}[\nu]|\mathcal{P}=p$ under the model. Note that $(\Theta_{p}\setminus\Theta_{*})[\nu]=\Theta_{p}\setminus\Theta_{*}$ by definition of $\Theta_{*}$, and thus for all $\nu\in S_{p}$ it holds that
$\mu^{\nu}_{p}(\Theta_{p}\setminus\Theta_{*})=\mu_{p}^{\text{id}}(\Theta_{p}\setminus\Theta_{*})=\Pi(\Theta_{p}\setminus\Theta_{*}\mid\mathcal{P}=p)=0.$ (B.1)
Lastly, Condition 9-(ii) means that $\lambda_{p}\ll\sum_{\nu\in S_{p}}\mu^{\nu}_{p}$, where $\ll$ denotes absolute continuity, and this together with (B.1) implies that $\lambda_{p}(\Theta_{p}\setminus\Theta_{*})=0$. But this is valid for all $p\in\mathbb{N}$, so we can conclude, as claimed, that $\lambda(\Theta\setminus\Theta_{*})=\sum_{p=1}^{\infty}\lambda_{p}(\Theta_{p}\setminus\Theta_{*})=0$. ∎
@@ C.1 Overview of data sets and comparison algorithms
A Gaussian process with a squared exponential kernel (also known as Gaussian kernel), namely $K_{4}(t,s)=e^{-(t-s)^{2}/2\nu^{2}}$, where $\nu=0.2$.
For the comparison algorithms themselves, we considered several frequentist methods which were selected among popular ones in FDA and machine learning in general. As specified in the main article, variable selection and dimensionality reduction methods are part of a pipeline followed by a standard multiple regression technique. In the linear regression case, we chose the following algorithms:
@@
Functional PLS regression through basis expansion, as in Aguilera et al. (2010).
@@
The main hyperparameters of all these algorithms were selected by 10-fold cross-validation, and for those that have a number of components to select, we set 10 as the maximum value so that comparison with our own methods are fair. In particular, regularization parameters are searched among 20 values in the logarithmic space $[10^{-4},10^{4}]$, the number of basis elements for cubic spline bases is in $\{4,5,\dots,10\}$, the number of basis elements for Fourier bases is one of $\{1,3,5,7,9\}$, and the number of neighbors in the KNN classifier is in $\{3,5,7,9,11,13\}$. Most algorithms have been taken from the libraries scikit-learn (Pedregosa et al., 2011) and scikit-fda (Ramos-Carreño et al., 2024), the first oriented to machine learning in general and the second to FDA in particular. However, some methods were not found in these packages and had to be implemented from scratch. This is the case of the FLDA, FPLS and APLS methods, which we coded following the corresponding articles.
@@ C.2 Simulations with non-Gaussian regressors
We performed an additional set of experiments in linear and logistic regression in which the regressors are not Gaussian processes (GPs), to see if our methods would hold up in this case. These experiments where run in the same conditions as those reported in the main article.
@@
We consider a “mixture” situation in which we combine regressors from two different GPs with equal probability and label them according to their origin. First, we consider a homoscedastic case to distinguish between a standard Brownian motion and a Brownian motion with a mean function that is zero until $t=0.5$, and then becomes $m(t)=0.75t$. Secondly, we consider a heteroscedastic case to distinguish between a standard Brownian motion and a Brownian motion with variance 2, that is, with kernel $K(t,s)=2\min\{t,s\}$.
Figure 8 shows that our classifiers perform better than most comparison algorithms in both cases. The differences are most notable in the homoscedastic case, and in the heteroscedastic case the overall accuracy is low. Incidentally, this heteroscedastic case of two zero-mean Brownian motions has a special interest, since it can be shown that the Bayes error is zero in the limit of dense monitoring (i.e. with an arbitrarily fine measurement grid), a manifestation of the “near-perfect” classification phenomenon analyzed for example in Torrecilla et al. (2020). Our results are in line with the empirical studies of this article, where the authors conclude that even though the asymptotic theoretical error is zero, most classification methods are suboptimal in practice (possibly due to the high collinearity of the data), with the notable exception of PCA+QDA.
@@ C.3 Analysis and validation of a model
We give an example through a series of visual representations of how one would analyze the outcome of our Bayesian methods. This is a preliminary step that comes before prediction; the idea is to validate the model and make sure that the resulting samples from the posterior are coherent and useful. For this illustration we consider a data set used in the experiments, for example the one with squared exponential GP regressors, and an underlying linear RKHS response given by (Figure 9)
@@
with $\varepsilon\sim\mathcal{N}(0,0.5)$. We run the sampler for $3000$ iterations and discard the first $2000$, with a $\text{Poisson}(3)$ prior truncated to $\{1,\dots,5\}$ for $p$.
@@
Next we proceed with the visual checks. We can look at the flat posterior distribution of all parameters for all values of $p$ and all the chains aggregated together (Figure 10), or visualize the traces of individual parameters for all values of $p$ (Figure 11). In addition, we can also look at the posterior of the multidimensional parameters for each $p$ in a sort of triangular configuration (Figure 12).
@@
Figure 12: Posterior distribution of $\tau_{p}^{*}$ for each $p$. The most frequent value is $p=3$, highlighted in red. In this run there were no samples with $p=1$ or $p=2$ after burn-in.
Looking at the traces is useful to check that the chains are mixing well and that they are correctly exploring the parameter space. We can do the same thing with the posterior values of $p$ (Figure 13). Moreover, we can visualize the tempered posterior distribution of $p$, that is, the posterior distribution of $p$ for each temperature (Figure 14).
@@
Lastly, we can perform a posterior predictive check (Figure 15). This is arguably the most useful visual test for prediction purposes, since we represent the posterior predictive distribution that will be used for inference and prediction. We can do it on the training data or directly on previously unseen regressors. If the sampling has been successful, the posterior predictive distribution should look like a tubular region around the observed data.
@@
C.4 Tables of experimental results
Here we present the tables corresponding to the empirical comparison studies in Appendix C.2 and in Section 4, which show the numerical values that were depicted there graphically. In each case the best and second-best results are shown in bold and blue, respectively.
@@ Appendix D Source code overview
  • The experiments folder contains plain text files with the numerical experimental results shown in Appendix C.4, as well as .csv files that facilitate working with them.
@@
The code for the experiments was executed with a random seed set to the value 2024 for reproducibility. We provide a script file launch.sh that illustrates a typical execution. Lastly, there are Jupyter notebooks that demonstrate the use of our methods in a more visual way. Inside these notebooks there is a step-by-step guide on how one might execute our algorithms, accompanied by many graphical representations, and offering the possibility of changing multiple parameters to experiment with the code. In addition, there is also a notebook that can be used to generate all the tables and figures of this document and the main article pertaining to the experimental results.

### v3

@@ B.2 Consistency and contraction rates via Schwartz’s theorem
Let $g$ be the density of $\Pi$ restricted to $\Theta_{0}$, and let $r_{0}>0$ be a radius such that $g(\theta)\geq g_{0}>0$ for all $\theta\in B(\theta_{0},r_{0})$. Then, for every $0<r\leq r_{0}$ we have
$\Pi(B(\theta_{0},r))=\int_{B(\theta_{0},r)}g\,d\lambda_{0}\geq g_{0}\lambda_{0}(B(\theta_{0},r))=g_{0}c_{d}r^{d},$
where $d=2p_{0}+2$, $\lambda_{0}$ is the Lebesgue measure on $\Theta_{0}$ and $c_{d}>0$ is a constant that depends only on the dimension of $\Theta_{0}$. ∎
Applying Lemma B.3 with $r=\varepsilon_{n}/\tilde{C}=Dn^{-\gamma}/\tilde{C}$ and taking logarithms in (B.15), we have, for large $n$,
$\log\Pi_{\mathcal{F}}(B_{2}(f_{0},\varepsilon_{n}))\geq\log c+(2p_{0}+2)\log\left(\frac{Dn^{-\gamma}}{\tilde{C}}\right)\geq\tilde{c}-2\gamma p_{0}\log n\geq-4\gamma p_{0}\log n,$
where $\tilde{c}>0$ is an irrelevant constant in the asymptotic regime. To finish the proof we want to find a constant $C>0$ such that $\log\Pi_{\mathcal{F}}(B_{2}(f_{0},\varepsilon_{n}))\geq-Cn\varepsilon_{n}^{2}=-CD^{2}n^{1-2\gamma}$, but $\log n$ grows much slower than $n^{1-2\gamma}$, so the inequality $4\gamma p_{0}\log n\geq-CD^{2}n^{1-2\gamma}$ holds for any choice of $C>0$ and sufficiently large $n$.
Proof of Corollary 12
We simply use the fact that, if $d$ is another distance on $\mathcal{F}$ with $d\leq\tilde{C}d_{H}$, then
$\Pi_{n}(\theta:d(f_{0},f_{\theta})>\varepsilon\mid\text{data})\leq\Pi_{n}(\theta:d_{H}(f_{0},f_{\theta})>\varepsilon/\tilde{C}\mid\text{data}).$
For the $L^{1}$-distance, by the Cauchy-Schwarz inequality we have
$d_{1}^{2}(f_{0},f_{\theta})$ $=\left(\int|f_{0}-f_{\theta}|\,d\rho\right)^{2}$
$=\left(\int\left|\sqrt{f_{0}}-\sqrt{f_{\theta}}\right|\left(\sqrt{f_{0}}+\sqrt{f_{\theta}}\right)d\rho\right)^{2}$
$\leq\int\left(\sqrt{f_{0}}-\sqrt{f_{\theta}}\right)^{2}d\rho\cdot\int\left(\sqrt{f_{0}}+\sqrt{f_{\theta}}\right)^{2}d\rho$
$\leq 8d^{2}_{H}(f_{0},f_{\theta}).$
In the last step we have used the alternative definition of the Hellinger distance, which can be expressed as $2d_{H}^{2}(f_{0},f_{\theta})=\int\left(\sqrt{f_{0}}-\sqrt{f_{\theta}}\right)^{2}d\rho$, the basic inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$, and the fact that both $f_{0}$ and $f_{\theta}$ integrate to $1$ over the whole space.
For the mean-variance discrepancy distance, recall the expression for the squared Hellinger distance between two normal densities, which can be formulated in this case as $d^{2}_{H}(f_{0},f_{\theta})=\int g_{0,\theta}(x)\,dQ_{X}(x)$, with
$g_{0,\theta}(x)=1-\sqrt{\frac{2\sigma_{0}\sigma}{\sigma_{0}^{2}+\sigma^{2}}}\exp\left\{-\frac{(\mu_{0}(x)-\mu_{\theta}(x))^{2}}{4(\sigma_{0}^{2}+\sigma^{2})}\right\}.$
Define $\Delta_{\mu}=\mu_{0}-\mu_{\theta}$, $\Delta_{\mu_{M}}=[\mu_{0}]_{M}-[\mu_{\theta}]_{M}$ and $S=\sqrt{\frac{2\sigma_{0}\sigma}{\sigma_{0}^{2}+\sigma^{2}}}$. On the one hand, using the trivial bounds $e^{-a}\leq 1$ and $S\leq 1$, we have
$g_{0,\theta}(x)\geq 1-S=\frac{1-S^{2}}{1+S}\geq\frac{1-S^{2}}{2}\geq C_{1}(\sigma_{0}^{2}-\sigma^{2})^{2},$
where the last inequality comes from the computations in (B.9). On the other hand, observe that $\Delta^{2}_{\mu}\geq\Delta^{2}_{\mu_{M}}$ and $\Delta^{2}_{\mu_{M}}\leq 4M^{2}$. Use these facts together with the bound $1-e^{-a}\geq\frac{a}{1+a}$ for $a\geq 0$ to get
$g_{0,\theta}(x)\geq 1-\exp\left\{-\frac{\Delta^{2}_{\mu_{M}}(x)}{4(\sigma_{0}^{2}+\sigma^{2})}\right\}\geq C_{2}\Delta^{2}_{\mu_{M}}(x).$
Bringing together both pieces of information, we can take $\tilde{C}=\min\{C_{1},C_{2}\}$ and write
$g_{0,\theta}(x)\geq\tilde{C}\left[\Delta^{2}_{\mu_{M}}(x)+(\sigma_{0}^{2}-\sigma^{2})^{2}\right],$
and integrating on both sides with respect to $Q_{X}$ yields
$d_{H}^{2}(f_{0},f_{\theta})\geq\tilde{C}\left[\|\Delta_{\mu_{M}}\|^{2}_{2,Q_{X}}+(\sigma_{0}^{2}-\sigma^{2})^{2}\right].$
To conclude, note that $\sqrt{a^{2}+b^{2}}\geq(1/\sqrt{2})(a+b)$ for $a,b\geq 0$, so we can take square roots on both sides and get the desired bound.
Examples of processes that satisfy Condition 8
Condition 8 implies that the process $X$ is “mean-square Lipschitz continuous”, i.e., there exists $L>0$ such that $\mathbb{E}[(X(t)-X(s))^{2}]\leq L(t-s)^{2}$ for all $t,s\in[0,1]$. Looking at the derivations we did, we can substitute the almost sure Lipschitzness of the trajectories by this new condition and arrive at the same conclusions. Moreover, note that this guarantees, via Kolmogorov’s continuity theorem, that the sample paths of $X$ (after an eventual modification) are continuous functions. Apart from somewhat trivial examples, such as a deterministic process or a linear random process, we can show that processes with a sufficiently smooth kernel satisfy our requirements.
Lemma B.4.
If the covariance function $K$ of the process $X$ is twice-continuously differentiable and the mean function $\mathbb{E}[X(t)]$ vanishes for all $t\in[0,1]$, then $X$ is “mean-square Lipschitz continuous”, that is, there exists $L>0$ such that $\mathbb{E}[(X(t)-X(s))^{2}]\leq L(t-s)^{2}$ for all $t,s\in[0,1]$.
@@
We know that $\mathbb{E}[(X(t)-X(s))^{2}]=K(t,t)-2K(t,s)+K(s,s)$, and a Taylor expansion around the point $(s,s)$ shows that
$K(t,t)-2K(t,s)+K(s,s)=\partial_{ts}K(s,s)(t-s)^{2}+o((t-s)^{2}).$
Since $\partial_{ts}K$ is continuous on the compact set $0,1]^{2}$, there exists a constant $L>0$ such that
$K(t,t)-2K(t,s)+K(s,s)\leq L(t-s)^{2},\quad\text{for }t\to s.$
Lastly, this bound extends to all $t,s\in[0,1]$ by continuity of $K$. ∎
Note that the above condition is by no means a necessary one. For example, given a zero-mean second-order stochastic process $\{Z(t):t\in[0,1]\}$, we consider the integrated process
$X(t)=\int_{0}^{t}Z(u)\,du,\quad t\in[0,1].$
If $K_{Z}(t,s)=\mathbb{E}[Z(t)Z(s)]$ is uniformly bounded above by a constant $L>0$ on $[0,1]$, we have
$\mathbb{E}\left[(X(t)-X(s))^{2}\right]=\int_{s}^{t}\int_{s}^{t}\mathbb{E}[Z(u)Z(v)]\,du\,dv\leq L(t-s)^{2}.$
For example, we might take $Z(t)$ to be a standard Brownian motion, an Ornstein-Uhlenbeck process, or, in general, any second-order process with a continuous covariance function. In any case, by virtue of Lemma B.4 we can always pre-process the trajectories so that they are smooth enough for the conditions to hold, considering for instance the convolution with a smooth kernel.
@@ C.1 Overview of data sets and comparison algorithms
A Gaussian process with a squared exponential kernel (also known as Gaussian kernel), namely $K_{4}(t,s)=e^{-(t-s)^{2}/2\xi^{2}}$, where $\xi=0.2$.
For the comparison algorithms themselves, we considered several frequentist methods which were selected among popular ones in FDA and machine learning in general. As specified earlier, variable selection and dimensionality reduction methods are part of a pipeline followed by a standard multiple regression technique. In the linear regression case, we chose the following algorithms:
@@
Functional PLS regression through basis expansion, implemented as in Aguilera et al. (2010).
@@
The main hyperparameters of all these algorithms were selected by 10-fold cross-validation, and for those that have a number of components to select, we set 10 as the maximum value so that comparison with our own methods are fair. In particular, regularization parameters are searched among 20 values in the logarithmic space $10^{-4},10^{4}]$, the number of basis elements for cubic spline bases is in $\{4,5,\dots,10\}$, the number of basis elements for Fourier bases is one of $\{1,3,5,7,9\}$, and the number of neighbors in the KNN classifier is in $\{3,5,7,9,11,13\}$. Most algorithms have been taken from the libraries scikit-learn (Pedregosa et al., 2011) and scikit-fda (Ramos-Carreño et al., 2024), the first oriented to machine learning in general and the second to FDA in particular. However, some methods were not found in these packages and had to be implemented from scratch. This is the case of the FLDA, FPLS and APLS methods, which we coded following the corresponding articles.
@@ C.2 Simulations with non-Gaussian regressors
We performed an additional set of experiments in linear and logistic regression in which the regressors are not Gaussian processes (GPs), to see if our methods would hold up in this case. These experiments where run in the same conditions as those reported in Section 4.
@@
We consider a “mixture” situation in which we combine regressors from two different GPs with equal probability and label them according to their origin. Firstly, we consider a homoscedastic case to distinguish between a standard Brownian motion and a Brownian motion with a mean function that is zero until $t=0.5$, and then becomes $m(t)=0.75t$. Secondly, we consider a heteroscedastic case to distinguish between a standard Brownian motion and a Brownian motion with variance 2, that is, with kernel $K(t,s)=2\min\{t,s\}$.
Figure 8 shows that our classifiers perform better than most comparison algorithms in both cases. The differences are most notable in the homoscedastic case, and in the heteroscedastic case the overall accuracy is low. Incidentally, this heteroscedastic case of two zero-mean Brownian motions has a special interest, since it can be shown that the Bayes error is zero in the limit of dense monitoring (i.e. with an arbitrarily fine measurement grid), a manifestation of the “near-perfect” classification phenomenon analyzed for example in Torrecilla et al. (2020). Our results are in line with the empirical findings of this article, where the authors conclude that even though the asymptotic theoretical error is zero, most classification methods are suboptimal in practice (possibly due to the high collinearity of the data), with the notable exception of PCA+QDA.
@@ C.3 Analysis and validation of a model
We give an example through a series of visual representations of how one would analyze the outcome of our Bayesian methods. This is a preliminary step that comes before prediction; the idea is to validate the model and make sure that the resulting samples from the posterior are coherent and useful. For this illustration we consider a data set used in the experiments, for example the one with squared exponential GP regressors and an underlying linear RKHS response given by
@@
with $\varepsilon\sim\mathcal{N}(0,0.5)$ (see Figure 9). We run the sampler for $3000$ iterations and discard the first $2000$, with a $\text{Poisson}(3)$ prior truncated to $\{1,\dots,5\}$ for $p$.
@@
Next we proceed with the visual checks. We can look at the flat posterior distribution of all parameters for all values of $p$ and all the chains aggregated together (Figure 10), or visualize the posterior of the multidimensional parameters for each $p$ in a sort of triangular configuration (Figure 11). In addition, we can also look at the traces of individual parameters for all values of $p$ (Figure 12).
@@
Figure 11: Posterior distribution of $\tau_{p}^{*}$ for each pp. The most frequent value is $p=3$, highlighted in red. In this run there were no samples with $p=1$ or $p=2$ after burn-in.
@@
Looking at the traces is useful to check that the chains are mixing well and that they are correctly exploring the parameter space. We can do the same thing with the values of $p$ (Figure 13). Moreover, we can visualize the tempered posterior distribution of $p$, that is, the posterior distribution of $p$ for each temperature (Figure 14).
@@
Lastly, we can perform a posterior predictive check (Figure 15). This is arguably the most useful test for prediction purposes, since we represent the posterior predictive distribution that will be used for inference and prediction. We can do it on the training data or directly on previously unseen regressors. If the sampling has been successful, the posterior predictive should look like a tubular region around the observed data.
@@
C.4 Execution times
In Figure 16 and Figure 17 we show the execution times of all the experiments in Appendix C.2 and Section 4.
Figure 16: Mean and standard error of execution times for all splits in the experiments with the functional linear model.
Figure 17: Mean and standard error of execution times for all splits in the experiments with the functional logistic model.
The execution times of the reference methods includes the duration of the cross-validation phase to select the best hyperparameters, a phase that our Bayesian methods lack by design. It is widely known that MCMC sampling tends to be slow, but as we can see the differences are manageable and our methods have reasonable run times for practical use, especially when taking into account the predictive improvement obtained in some cases.
In addition, it is possible that we ran some MCMC chains for more steps than necessary, because we wanted to perform all experiments in the same generic configuration. In real-world scenarios one would pay closer attention to convergence metrics and try to stop the sampling earlier. Moreover, some splits may be artificially long because of the queue system in the cluster used to run the experiments in parallel.
C.5 Tables of experimental results
Here we present the tables corresponding to the empirical comparison studies in Appendix C.2 and Section 4, which show the numerical values that were depicted there graphically. In each case the best and second-best results are shown in bold and italicized blue, respectively.
@@ Appendix D Source code overview
  • The results folder contains plain text files with the execution times and the numerical results shown in Appendix C.4 and Appendix C.5, as well as .csv files that facilitate working with them.
@@
The code for the experiments was executed with a random seed set to the value 2024 for reproducibility. We provide a script file launch.sh that illustrates a typical execution. Lastly, there are Jupyter notebooks that demonstrate the use of our methods in a more visual way. Inside these notebooks there is a step-by-step guide on how one might execute our algorithms, accompanied by many graphical representations, and offering the possibility of changing multiple parameters to experiment with the code. In addition, there is also a notebook that can be used to generate all the tables and figures of this document pertaining to the experimental results.